I am having serious dificculty on finding out speed of the express train and the local train.
I know that I double the speed of the express train, by the speed of the local one, but what step do I need to take to find out the speed of the local train?
Question:
An express and local train leave Gray's lake at 3 pm and head for Chicago 50 miles away. The express travels twice as fast as the local and arrives 1 hour ahead of it. Find the speed of each train.

Answer by ptaylor(2198)


Distance(d) equals Rate(r) times Time(t) or d=rt; r=d/t and t=d/r
Let r=rate of the local train
Then 2r=rate of the express
Time of local train=50/r
Time of express train=50/2r
Now we are basically told the following:
(50/r)-1=50/2r multiply each side by 2r
100-2r=50 subtract 100 from each side
100-100-2r=50-100 collect like terms
-2r=-50 divide each side by -2
r=25 mph-------------speed of local
2r=2*25=50 mph--------------speed of express
CK
(50/25)-1=50/50
1=1

Hope this helps---ptaylor